Session: Nov 2000 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

\(Other\) (M.Bondi at Fri Nov 17 07:30:20 MET 2000) Success

Medicina (M.Bondi at Fri Nov 17 07:59:09 MET 2000) Success

Torun (K.Borkowski at Fri Nov 17 09:11:53 MET 2000) Success
No known problems.

Onsala (A. Polatidis at Fri Nov 17 13:01:39 MET 2000) Minor failures
We have some problems with the recording quality, very high parity errors, some in $$$$. We are investigating the problem. The Formatter jumped by 1 sec before the beggining of the experiment. The operator resynched the formatter at 0415 UT

Effelsberg (A. Kraus at Sun Nov 19 16:34:14 MET 2000) Minor failures
Tsys measurements between UT0430-0451 bad.

Jodrell_Bank (Alastair Gunn at Mon Nov 20 15:15:25 MET 2000) Success
No known problems.

Noto (C.Trigilio at Mon Nov 20 16:38:51 MET 2000) Success

Westerbork (Tony Foley at Wed Nov 22 09:13:56 MET 2000) Minor failures
Major problems with the array control software in the period just before startup. RT0 out of array (so 13 telescopes in). At the end (from about 05:53UT) overflow on BBC1. Parity check during experiment suggested poor recording, but I played some back afterwards and it looked OK, but shifted about 16microns

Westerbork (Tony FOley at Wed Nov 22 11:10:38 MET 2000) Minor failures
The scan 322-0623 has source outside HA limits, so not observed